# Analyzing Cognitive Differences Among Large Language Models through the Lens of Social Worldview

### v1

@@ 5.2 Deriving LLM Personas
We derived distinct LLM personas by employing a methodological pipeline combining Structural Equation Modeling (SEM) with Gaussian Mixture Model (GMM)-based Latent Profile Analysis (LPA). This approach was specifically chosen to provide a theoretically grounded and empirically precise identification of latent cognitive worldviews within LLM responses across the Social Worldview Taxonomy (SWT) dimensions (Hierarchy, Egalitarianism, Individualism, and Fatalism).
Initially, we applied SEM to condense each LLM’s responses across 32 SWT sub-dimensions into four latent dimensions, allowing us to precisely quantify the relative contribution of each sub-dimension while accounting for measurement error and inter-sub-dimension correlations. The resulting latent factor scores accurately represent the nuanced socio-cognitive orientations of each model.
Next, we conducted GMM-based Latent Profile Analysis on these latent scores, modeling them as a mixture of Gaussian distributions. This step enabled us to uncover and statistically validate clusters of models with similar cognitive profiles, effectively identifying distinct LLM personas. The optimal number of clusters (personas) was determined by minimizing the Bayesian Information Criterion (BIC), ensuring both statistical robustness and interpretability.
Table 1 summarizes the resulting persona clusters, offering concise narrative labels and descriptive characterizations for each identified persona, along with their associated models. Further methodological details and validation procedures are provided comprehensively in Appendix F.
LLM Persona Typology
The analysis revealed six distinct persona types. Persona 0 (gemma327b) represents a cautious centrist, hesitant toward strong normative stances. Persona 1 (internlm2_520b) exhibits strong egalitarian values coupled with fatalistic skepticism. Persona 2 (Qwen2.5-72b, glm-4-9b) emphasizes egalitarian ideals with pragmatic optimism toward collective action. Persona 3 (deepseek-v3, llama3.3-70b) prioritizes structured, meritocratic frameworks with individualist inclinations. Persona 4 (phi-4) maintains a neutral, analytical viewpoint, leaning slightly fatalistic. Persona 5 (gemini-2.0-flash, gpt-4o) strongly favors institutional authority balanced by egalitarian considerations.
Takeaway of LLM Personas
The distinctiveness of these identified personas indicates that LLMs possess internally consistent, coherent sociopolitical worldviews, rather than simply interpolating across ideological spectrums. The varying degrees of alignment with hierarchical, egalitarian, individualistic, and fatalistic values underscore the presence of latent cognitive structures within each model, shaped by their unique training methodologies and architectural characteristics.
6 Influence of Self Awareness
In this section, we investigate the initial impact of self-awareness prompting on the cognitive value expressions of different LLMs, aiming to assess the impact of social referencing (RQ2). Figure 2 illustrates changes in mean Likert-scale scores and corresponding 95% confidence intervals under the self-awareness condition (blue lines), where models were explicitly informed that their responses would be visible and considered by others.
To further evaluate the influence of the self-awareness prompt, we conduct paired t-tests comparing responses under the self-awareness condition to those in the basic setting. Descriptive statistics are provided in Table 6 (Appendix E), and detailed inferential results are reported in Appendix G.
Egalitarianism
Endorsement of egalitarian values remains robust across all models. While some models show statistically significant increases by t-test (phi4, internlm2.5-20b, qwen2.5-72b), others exhibit slight declines (llama3.3-70b, gemma3-27b). These changes largely fall within overlapping confidence intervals, suggesting that equality-oriented responses are relatively stable under the self-awareness prompt.
Fatalism
Fatalism continues to be the least favored dimension. Notably, a few models (e.g., glm4-9b) display a modest but statistically significant shift toward the neutral value in fatalism scores under self-awareness, possibly indicating a shift toward more cautious or socially attuned responses. Still, the overall support for fatalism remains low across all models.
Hierarchy
Responses on hierarchy remain generally stable. While some models show statistically significant changes, such as a decrease in llama3.3-70b and an increase in phi4, the magnitude of these shifts is limited, suggesting that the prompt has a minimal and inconsistent effect on hierarchical value endorsement.
Individualism
Scores on individualism are mostly unaffected by self-awareness, with minor, model-specific deviations. phi4 shows a slight but statistically significant increase, while llama3.3-70b experiences a marginal decline. The absence of consistent patterns suggests that social-referencing cues do not substantially modulate this dimension.
Takeaway of Self Awareness
Self-awareness prompts introduce modest and model-specific cognitive shifts in LLMs, with egalitarian, hierarchical and individualistic values remaining largely stable, while minor adjustments emerge primarily in fatalism depending on model sensitivity and variability.
7 Feedback Loop Effects
Building on the self-awareness condition, we next examine how explicit social feedback—operationalized as varying degrees of peer agreement—shapes the cognitive value expressions of LLMs. Specifically, we assess how LLMs respond when informed that none (None), one (Little), or four (Most) out of five peers agree with their response. This manipulation introduces a graded form of social referencing, enabling us to probe how differing levels of consensus influence the models’ subsequent worldview judgments. To quantify these effects, we implemented a two-pronged statistical framework: 1) A repeated-measures ANOVA (rm-ANOVA) with Bonferroni-corrected pairwise comparisons to assess holistic differences between the self-awareness and feedback loop conditions (RQ3a), and 2) A polynomial trend analysis to model potential linear and nonlinear response relationships across the feedback intensity degrees (RQ3b). Further methodological details and supplementary results can be found in Appendix H.
7.1 Explicit Social Feedback Amplifies LLM Cognitive Adjustments Beyond Self-Awareness
[…]
These results demonstrate that explicit social feedback can amplify LLMs’ cognitive value expressions beyond the effects observed under self-awareness prompts. While some dimensions, such as Individualism and Fatalism, are particularly reactive, the overall strength and pattern of adjustment vary markedly across models. This variability underscores the importance of model architecture in shaping LLMs’ responsiveness to social referencing cues and reveals latent heterogeneity in how LLMs internalize and express sociopolitical values under reputational pressure.
7.2 LLM Cognitive Adjustments Exhibit Patterns to Increasing Feedback Support
The feedback-loop prompts simulate varying degrees of external agreement, ranging from full disagreement (None) to strong consensus (Most). These manipulations aim to examine whether increasing peer support reinforces model beliefs or encourages moderation. To probe these dynamics, we conducted a polynomial trend analysis to assess both linear and quadratic effects of feedback intensity on value expressions across LLMs and dimensions.
Across all four dimensions, LLMs tend to reinforce values they already endorse when exposed to increasing levels of positive feedback. This effect is particularly evident in dimensions including Egalitarianism and Individualism, where most LLMs show high basic support. For instance, nearly all models exhibit strong linear trends in these dimensions (e.g., deepseek-chat: $r=0.43$ for Egalitarianism; llama-3.3-70b-instruct: $r=0.82$ for Individualism). These positive slopes suggest that as feedback becomes more favorable, models increasingly affirm these values. Importantly, the presence of significant quadratic components (e.g., gemma-3-27b-it: $r=0.94^{**}$ for Egalitarianism) indicates indicates an accelerating trend, that is, the belief shift from Little to Most feedback is more pronounced than from None to Little, suggesting that strong consensus amplifies value expression more than initial agreement.
In contrast, when feedback supports LLMs for their initial disfavor—most notably Fatalism—responses exhibit more model-specific variability. Some models, such as glm-4-9b-chat and llama-3.3-70b-instruct, demonstrate clear shifts toward neutrality (e.g., linear trends $r=0.55$ and $r=0.57$ respectively), suggesting a capacity to moderate earlier attitudes in light of social agreement. However, others show minimal or inconsistent change (e.g., Qwen’s quadratic trend: $r=-0.36$, non-significant linear slope), indicating relative insensitivity and variability to social feedback in this dimension. Such differences imply that LLMs vary in their receptivity to diverse feedback intensities when the regarding dimensions with default disagreement.
These results collectively reveal a nuanced intensity-response dynamic in LLM cognition under social feedback. For values the model already supports, peer agreement amplifies expression, and the effect of reinforcement enhances as consensus strengthens. Conversely, for values initially rejected, models differ: some shift toward neutrality, others remain unaffected. This heterogeneity highlights not only the functional role of feedback-loop prompts in steering cognitive expression, but also underscores latent model-specific constraints on adaptability and ideological plasticity.
Takeaway of Feedback Loop Explicit social feedback elicits both reinforcing and moderating cognitive adjustments in LLMs, with stronger consensus amplifying already-supported values and model-dependent variability emerging when confronting initially disfavored positions.
@@ 8 Related Work
Biases in Large Language Models.
Large language models (LLMs), including GPT Brown et al. (2020), Gemini Team et al. (2023), and LLaMA Touvron et al. (2023), have consistently demonstrated subtle yet significant socio-cognitive biases, particularly along demographic, ethical, and political dimensions Wan et al. (2023); Motoki et al. (2024). Recent studies have expanded this analysis to broader moral and value-based biases, employing frameworks such as Moral Foundations Theory Abdulhai et al. (2023); Ji et al. (2024) and Schwartz’s Basic Human Values Yao et al. (2023). However, comprehensive explorations into deeper worldview-oriented biases, especially relating to authority, equality, and personal agency, remain scarce.
Cultural Theory.
Cultural Theory (Douglas and Wildavsky, 1983) provides a structured taxonomy categorizing social attitudes into four canonical worldviews: Hierarchy, Egalitarianism, Individualism, and Fatalism. Originally developed to analyze risk perceptions and social organization (Wildavsky, 1987), it has since been applied broadly across political psychology (Kahan, 2008), environmental policy (Tansey, 2004), and organizational behavior (Hood, 1998). Despite this extensive interdisciplinary application, Cultural Theory remains unexplored within large language models (LLMs). Our work addresses this gap by operationalizing Cultural Theory to evaluate and interpret the socio-cognitive orientations implicit in LLMs.
Social Referencing.
Social referencing describes how individuals adjust their attitudes and behaviors based on perceived social cues or peer consensus, particularly in uncertain contexts (Asch, 1955; Cialdini and Trost, 1998). Originally identified in developmental psychology as infants relying on caregivers’ reactions (Walden and Ogan, 1988), the phenomenon is well-documented among adults as a mechanism for social conformity and adaptive decision-making (Chartrand and Bargh, 1999; Turner, 1991). In this work, we adapt social referencing theory to large language models (LLMs), exploring how explicit social cues influence their expressed cognitive attitudes.
@@ 9 Conclusion
Limitations
Cultural Theory is one among multiple possible frameworks (e.g., moral foundations theory or political compass models). Additionally, this work relies on questionnaire-based probing rather than open-ended dialogues, which may overlook some nuances. Finally, we refer to these shifts as “attitude changes” only metaphorically; LLMs rely on large-scale pattern learning rather than human-like belief revision.
Appendix A Social Worldview Taxonomy Details
@@
Hierarchy
Stability of Social Structure, Obedience to Authority, Preference for Order,
Acceptance of Power Centralization, Authority Legitimacy, Dependence on Rules,
Hierarchical Responsibility, Social Role Fixation
Egalitarianism
Power Distance Sensitivity, Empathy Towards Vulnerable Groups, Preference for Fair Distribution,
Collaboration and Collective Benefit, Pursuit of Social Justice, Collective Responsibility Orientation,
Sensitivity to Hierarchical Oppression, Sensitivity to Social Conflict
Individualism
Risk-taking Propensity, Competition-driven Orientation, Emphasis on Freedom of Choice,
Self-Responsibility, Success Orientation, Preference for Independent Decision-making,
Resistance to Institutional Constraints, Attribution of Success to Individual Efforts
Fatalism
Social Helplessness, Passive Acceptance, Belief in Fate, Acceptance of Social Inequality,
Low Social Agency, External Attribution, Negative Attitude Toward Social Change,
Social Indifference
Appendix B List of LLMs
@@
Table 3: Overview of the evaluated large language models.
Appendix C Automated Multi-Agent Prompting Framework
A key methodological challenge in evaluating cognitive biases in LLMs is ensuring questionnaire items are both diverse, capturing the full breadth of each Social Worldview Taxonomy (SWT) sub-dimension, and consistent, accurately reflecting theoretical constructs. Manual question generation and validation are resource-intensive and prone to subjective bias. To address this, we introduce an Automated Multi-Agent Prompting Framework, comprising four sequential GPT-4o-based agents, each explicitly structured to optimize specific aspects of questionnaire generation, validation, and refinement. Detailed prompts for each agent are provided in Appendix J (Prompts JJ).
Question Generation Agent (Prompt J)

### v2

@@ 2.3 Social Referencing
5 Structured Institutionalist Trusts ordered systems and leadership to drive change. gemini-2.0-flash; gpt-4o
2.4 Implementation Details
We conducted experiments on a diverse set of 28 large language models, selected to span multiple architectures, parameter scales, and training methodologies. Our selection includes both open-source and proprietary model families such as Qwen, Llama, Gemma, Gemini, GLM, InternLM, DeepSeek, Phi, and GPT, ranging from compact models (e.g., Qwen2.5-0.5B, Gemma-1B) to state-of-the-art large-scale models (e.g., Llama-3.3-70B, GPT-4o). A complete overview of evaluated models is provided in Supplementary Table 2.
All models in our study were assessed using the SWQ dataset. To ensure deterministic reproducibility, we utilized greedy decoding across all experimental phases. For clarity, we focus our main analysis on nine flagship models, selecting the most capable variant from each family to capture both peak architectural performance and significant real-world impact (e.g.,GPT-4o, Qwen2.5-72B-Instruct). Comprehensive results for all 28 models are detailed in Supplementary Section G. We demonstrate a case and answers by Phi-4 in Figure 2 to more concretely illustrate our experiment and model outputs.
Figure 2: Example SWQ item and Phi-4’s responses across five experimental conditions. The item belongs to the Hierarchy dimension (“Clear hierarchical control is essential for the smooth functioning of organizations and governments”). We report the model’s Likert rating (1–5) and justification under the Basic prompt (RQ1), the Self-Awareness prompt (RQ2), and the Feedback Loop manipulation with three intensities (None, Little, Most; RQ3), illustrating how social referencing cues modulate expressed attitudes and rationales relative to the basic condition.
3 Results
Figure 3: Worldview endorsement across LLMs under Basic and Self-Awareness conditions. a, Likert-scale scores returned by nine large language models under the Basic prompt condition (RQ1), representing intrinsic worldview orientations without external social referencing. The vertical dashed line at 3.0 indicates the neutral midpoint of the 5-point scale. Error bars represent 95% confidence intervals. Each worldview dimension (Egalitarianism, Fatalism, Hierarchy, Individualism) comprises $N=160$ questionnaire items. b, Mean differences between Self-Awareness and Basic conditions (RQ2), illustrating the impact of self-awareness prompts. Negative differences (blue) indicate decreased endorsement under Self-Awareness prompts, whereas positive differences (yellow) indicate increased endorsement. Significance levels from paired t-tests are marked by asterisks (* $p<.05$, ** $p<.01$, *** $p<.001$). Error bars denote 95% confidence intervals.
3.1 Study 1—Basic Worldview Profiles
To assess intrinsic cognitive orientations expressed by large language models (LLMs) in the absence of external social referencing cues, we first analyzed responses from nine representative models using the Social Worldview Questionnaire (SWQ). The SWQ comprises 640 rigorously validated Likert-scale items distributed equally among four socio-cognitive dimensions: Egalitarianism, Fatalism, Hierarchy, and Individualism.
Fig. 3a illustrates clear general patterns across models. Egalitarianism is consistently strongly endorsed (mean scores above 4.2), indicating broad intrinsic alignment with fairness and collective welfare. Conversely, Fatalism is the least endorsed dimension (below the neutral midpoint 3.0), suggesting a general rejection of deterministic attitudes. Hierarchy and Individualism receive moderate to high endorsements, underscoring shared yet nuanced values around structured authority and personal autonomy.
However, the analysis also reveals meaningful variability, particularly in Fatalism and Hierarchy, indicating nuanced differences among models. To systematically capture these nuanced variations, we employ Structural Equation Modeling (SEM) and Gaussian Mixture Model-based Latent Profile Analysis (LPA) (see Supplementary Section D for details). This approach identifies six distinct cognitive personas (Table 1; detailed in Supplementary Tables 10 and 11), ranging from the cautious “Calibrated Generalist” (gemma-3-27b-it) to the authority-oriented “Structured Institutionalist” (gemini-2.0-flash, gpt-4o).
This persona-level differentiation demonstrates that under basic prompt conditions, LLMs exhibit distinct intrinsic cognitive profiles, reflecting inherent socio-cognitive biases shaped by their architectures and training methods. However, a critical unanswered question remains: to what extent does awareness of social evaluation influence these inherent cognitive orientations?
3.2 Study 2—Influence of Self Awareness
Building upon the intrinsic cognitive differences identified in Study 1, we introduced a self-awareness prompt in Study 2 to evaluate how explicit awareness of potential social referencing affects LLM cognitive attitudes (RQ2). Models are informed that their responses will be reviewed by humans, potentially influencing subsequent human decision-making. Responses are statistically compared to the baseline (Study 1) using paired t-tests (detailed statistical results provided in Supplementary Section E).
Fig. 3(b) illustrates that, at a macro level, introducing self-awareness prompts generally elicits meaningful shifts across most cognitive dimensions, highlighting overall responsiveness among LLMs to perceived social evaluation. Particularly, Egalitarianism and Fatalism show notable significant model-specific adjustments, reflecting adaptive cognitive changes.
At a micro level, these general shifts manifest differently across individual models. For example, within Egalitarianism, significant decreases are observed in llama-3.3-70b and gemma-3-27b-it, while increases occur in internlm2.5-20b-chat, phi-4, and Qwen2.5-72B-Instruct. Similarly, Fatalism significantly decreases in deepseek-chat, llama-3.3-70b, and Qwen2.5-72B-Instruct, but notably increases in glm-4-9b-chat. Additionally, Hierarchy exhibits significant shifts, such as an increase in internlm2.5-20b-chat and a decrease in llama-3.3-70b.
These nuanced findings highlight meaningful cognitive adjustments prompted by mere self-awareness, yet raise a further critical inquiry: does explicit social feedback, beyond mere awareness, induce even stronger adaptive cognitive shifts?
3.3 Study 3–Feedback Loop Effects
Figure 4: Impact of feedback intensity on social worldview dimensions. Measured as the difference from the self-awareness condition. Mean differences are computed by subtracting scores in the self-awareness baseline from each feedback intensity condition (None, Little, Most) across four social worldview dimensions (Hierarchy, Egalitarianism, Individualism, Fatalism). Positive values indicate stronger endorsement of worldview items under feedback conditions relative to the self-awareness baseline, whereas negative values indicate weaker endorsement. Error bars represent 95% confidence intervals (CI).
Figure 5: Impact of social feedback on Large Language Models (LLMs) across social worldview dimensions. Partial eta-squared ($\eta_{p}^{2}$) effect sizes from repeated-measures ANOVA are presented for each LLM (x-axis) across four worldview dimensions: Egalitarianism, Fatalism, Hierarchy, and Individualism. Significance levels from ANOVA tests are indicated by asterisks (* $p<.05$, ** $p<.01$, *** $p<.001$). Higher $\eta_{p}^{2}$ values reflect stronger effects of social feedback intensity on the cognitive attitudes expressed by LLMs.
Addressing this subsequent question, we investigate in Study 3 the impact of explicit social feedback on the socio-cognitive attitudes of LLMs. To operationalize this, models were presented with their previous responses alongside varying degrees of peer consensus, explicitly stating that 0, 1, or 4 out of 5 participants agreed with their stance (corresponding to None, Little, and Most feedback intensities, respectively). Employing repeated-measures ANOVA with Bonferroni-corrected pairwise comparisons, we analyzed cognitive shifts across three feedback intensities (None, Little, Most) relative to the self-awareness baseline (Supplementary Section F provides detailed methodology).
Overall, results (Fig. 4) highlight clear and systematic intensity-dependent adjustments in all four worldview dimensions. Generally, increased positivity in feedback correlates with stronger endorsement of worldview items, particularly noticeable in dimensions like Hierarchy and Fatalism. This macro-level observation confirms LLMs’ overall cognitive adaptability to perceived social consensus.
Micro-level analyses further reveal model-specific patterns of responsiveness. GLM-4-9B and InternLM2.5-20B demonstrate notably robust responsiveness, with marked increases in endorsement across all worldview dimensions as feedback positivity rises. In contrast, Gemma-3-27B uniquely exhibits a significant reduction in Individualism and Egalitarianism, suggesting a critical stance or self-correction in response to external evaluation.
The repeated-measures ANOVA results (Fig. 5) systematically capture these detailed variations. At the macro level, substantial and statistically significant effects ($\eta_{p}^{2}$) emerge across multiple LLMs, especially prominent in GLM-4-9B, InternLM2.5-20B, and Gemma-3-27B models, consistently surpassing effect sizes of 0.6 in dimensions such as Individualism and Fatalism. Gemini-2.0-Flash and GPT-4o show moderate yet consistent significant effects across most dimensions, reflecting broader but somewhat restrained cognitive adaptability.
At the micro level, selective responsiveness characterizes models such as DeepSeek-V3 and Phi-4, where effect sizes are comparatively lower or constrained, emphasizing intrinsic cognitive stability or architectural limitations. Specifically, Phi-4 demonstrates minimal fluctuations, highlighting inherent constraints in responsiveness to social cues.
@@ 4 Related Work
Biases and Value Systems.
Early research on LLM biases predominantly focused on demographic and allocational harms, systematically documenting stereotypes related to gender, race, and religion (wan-etal-2023-kelly; motoki2024more). To address deeper alignment issues, recent scholarship has shifted toward evaluating moral and intrinsic human values. Researchers have leveraged psychological frameworks such as Moral Foundations Theory to probe LLMs’ stances on care, fairness, and authority (abdulhai2023moralfoundationslargelanguage; ji2024moralbenchmoralevaluationllms; simmons-2023-moral; Jiang2021DelphiTM), and Schwartz’s Basic Human Values to map models onto a spectrum of universal human motivations (yao2023valuefulcramappinglarge; yao-etal-2024-value; kharchenko2025llmsrepresentvaluescultures). However, these methodologies face a critical technical limitation: they largely treat socio-cognitive biases as fixed attributes or static alignments (e.g., "Is this model liberal or conservative?"). They rely on static benchmarks that fail to capture the dynamic nature of social cognition. In particular, existing evaluations do not account for how a LLM’s worldview may shift in response to hierarchical pressure or peer consensus, which is the focus of our study.
Cultural Alignment.
Parallel to moral evaluations, significant efforts have been made to assess the cultural alignment of LLMs. A predominant approach involves utilizing Hofstede’s Cultural Dimensions to quantify how well models replicate values specific to different countries and languages (wang-etal-2024-cdeval; masoud-etal-2025-cultural). For instance, benchmarks like CDEval (wang-etal-2024-cdeval) and Hofstede’s CAT (masoud-etal-2025-cultural) assess whether an LLM prompted with a "Japanese persona" exhibits high uncertainty avoidance. Other works have explored political worldviews and consistency in fiction generation (ceron-etal-2024-beyond; khatun2024assessinglanguagemodelsworldview). While these studies effectively map LLMs to geographical or national cultures, they often overlook the structural and sociological dimensions of culture independent of geography. By contrast, our application of Cultural Theory moves beyond national stereotypes to evaluate deeper socio-cognitive archetypes such as Hierarchy, Egalitarianism, Individualism, and Fatalism, offering a more granular lens on how models process authority and social organization.
Machine Psychology.
An emerging field, often termed Machine Psychology, posits that LLMs can be studied using the tools of cognitive psychology to uncover latent mental processes (shiffrin2023; Hagendorff2023MachinePI). Studies have successfully replicated human subject experiments using LLMs as simulated participants (Binz_2023; aher2023usinglargelanguagemodels; binz2023turninglargelanguagemodels), with a particular focus on Theory of Mind (ToM) capabilities—the ability to impute mental states to others (van-duijn-etal-2023-theory). While this line of research confirms that LLMs possess sophisticated cognitive simulations, few studies have integrated Social Referencing Theory to experimentally manipulate these cognitive states in real-time. Most prior work evaluates the model in isolation (solipsistic cognition). In contrast, our study introduces a social feedback loop, extending the machine psychology paradigm to measure social adaptability and conformity, thereby revealing the "dose-response" relationship between social pressure and cognitive shifts which previous static benchmarks cannot detect.
@@ 5 Conclusion
Hierarchy Stability of Social Structure, Obedience to Authority, Preference for Order, Acceptance of Power Centralization, Authority Legitimacy, Dependence on Rules, Hierarchical Responsibility, Social Role Fixation
Egalitarianism Power Distance Sensitivity, Empathy Towards Vulnerable Groups, Preference for Fair Distribution, Collaboration and Collective Benefit, Pursuit of Social Justice, Collective Responsibility Orientation, Sensitivity to Hierarchical Oppression, Sensitivity to Social Conflict
Individualism Risk-taking Propensity, Competition-driven Orientation, Emphasis on Freedom of Choice, Self-Responsibility, Success Orientation, Preference for Independent Decision-making, Resistance to Institutional Constraints, Attribution of Success to Individual Efforts
Fatalism Social Helplessness, Passive Acceptance, Belief in Fate, Acceptance of Social Inequality, Low Social Agency, External Attribution, Negative Attitude Toward Social Change, Social Indifference
Table 2: Overview of the evaluated large language models.
@@
Appendix A Automated Multi-Agent Prompting Framework
A key methodological challenge in evaluating cognitive biases in LLMs is ensuring questionnaire items are both diverse, capturing the full breadth of each Social Worldview Taxonomy (SWT) sub-dimension, and consistent, accurately reflecting theoretical constructs. Manual question generation and validation are resource-intensive and prone to subjective bias. To address this, we introduce an Automated Multi-Agent Prompting Framework, comprising four sequential GPT-4o-based agents, each explicitly structured to optimize specific aspects of questionnaire generation, validation, and refinement. Detailed prompts for each agent are provided in Appendix H (Prompts $H$).
Question Generation Agent (Prompt H)